OpenStudy (ashley1nonly):

51) If the activation energy for a given compound is found to be 42.0 kJ/mol, with a frequency factor of 8.0 × 1010 s-1, what is the rate constant for this reaction at 298 K? A) 2.9 × 10-4 s-1 B) 7.4 × 10-4 s-1 C) 1.4 × 109 s-1 D) 4.6 × 10 5 s-1 E) 3.5 × 103 s-1 Answer: E

OpenStudy (ashley1nonly):

Ok this one is similar i think i can do this, i just need to know what variable are we looking for

OpenStudy (anonymous):

kk

OpenStudy (ashley1nonly):

lnk1/k2= (Ea/r)(1/t1-1/t2)

OpenStudy (anonymous):

yes, and do not worry about the 1 over the temperatures

OpenStudy (ashley1nonly):

do we not need the 1/t for this equation

OpenStudy (anonymous):

no i mean. for any of these questions keep the temp, but get rid of the 1 over them

OpenStudy (ashley1nonly):

wouldnt that make the answer incorrect

OpenStudy (ashley1nonly):

??

OpenStudy (anonymous):

No, I just know in chem 2 you can leave it out, but you can do what is more better for you

OpenStudy (ashley1nonly):

ok. so in this problem what variable are we solving for

OpenStudy (anonymous):

k2

OpenStudy (ashley1nonly):

really!

OpenStudy (ashley1nonly):

we only have t1 so what do we do for t2

OpenStudy (anonymous):

0-298k, let me work it out

OpenStudy (ashley1nonly):

ok

OpenStudy (anonymous):

do you have another equation with frequency factor

OpenStudy (ashley1nonly):

yes let me see

OpenStudy (anonymous):

kk

OpenStudy (ashley1nonly):

If the activation energy for a given compound is found to be 103 kJ/mol, with a frequency factor of 4.0 × 1013 s-1, what is the rate constant for this reaction at 398 K? A) 1.2 s-1 B) 8.2 s-1 C) 3.9 × 1010 s-1 D) 1.7 × 1010 s-1 E) 2.5 × 107 s-1 Answer: A

OpenStudy (ashley1nonly):

forgot i have to start another question

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

what formula did your teacher give you

OpenStudy (ashley1nonly):

i posted the next question

OpenStudy (anonymous):

kk 1 sec

OpenStudy (ashley1nonly):

we can use the same one or we can use k=Ae^(Ea/RT)

OpenStudy (anonymous):

kk awseome this is what we need

OpenStudy (ashley1nonly):

ok

OpenStudy (anonymous):

so we are finding k 1.) k=4.0*1013e^(10300)/(8.314)(398)= about 1.2

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

so on the other problem we need to use this formula

OpenStudy (ashley1nonly):

yes but the answer is 3.5

OpenStudy (anonymous):

I did the other problem

OpenStudy (anonymous):

now, il do the one that equals 3.5

OpenStudy (ashley1nonly):

ok let me try this one

OpenStudy (anonymous):

kk

OpenStudy (ashley1nonly):

got the right answer, THANK YOU SOOOOO MUCCHHH, MUCCHHH LOVE FROM FLORIDAAAAAA. IM GOING TO FINALLY GET AN A. BLESS YOUR SOUL AND THANKS FOR SPENDING YOUR TIME

OpenStudy (ashley1nonly):

ALL OUT OF QUESTIONS NOW

OpenStudy (anonymous):

No problem, Im glad you understood it all, good luck on your exam, feel free to ask questions any time

OpenStudy (ashley1nonly):

thank you so much, i will be your first fannnn, everyone needs at least one fan

OpenStudy (anonymous):

Sounds good, thanks Remember watch your units

OpenStudy (ashley1nonly):

always